Find how many qrts of 6% butterfat milk and 2% butterfat milk should be mixed to yield 120 qrts of 5% butterfat milk
I think it is 90 of the 6% and 30 of the 2%
Is this correct

Answer by Alan3354(69443)

Find how many qrts of 6% butterfat milk and 2% butterfat milk should be mixed to yield 120 qrts of 5% butterfat milk
I think it is 90 of the 6% and 30 of the 2%
Is this correct
----------------
90*0.06 + 30*0.02 = 5.4 + 0.6 = 6
6/120 = 0.05 = 5%
It's correct